OpenStudy (anonymous):

what is the integral of sec^5 x dx

OpenStudy (adamaero):

You need u-substitution for this one. When secx is to the odd power, you should automatically think: make u = ?

OpenStudy (anonymous):

how can i do that? I have conflicts aboutthis opic

OpenStudy (adamaero):

This info should be in the beginning of whatever chapter you're on or a past one if you've already covered it. You'll learn more if you try to find it on your own.

OpenStudy (anonymous):

someone please help me.

OpenStudy (anonymous):

\[\sec ^5x=\sec ^4x \sec x=\left( \sec ^2x \right)^2\sec x=\left( 1+\tan ^2x \right)^2\sec x\] now you can solve.

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

yw